- Press Release
- Mar 24, 2023
Mars Global Surveyor Status Report 04-05-2000
Wednesday, April 5, 2000 (DOY 089/19:00:00 to DOY 096/19:00:00 UTC)
Launch / Days since Launch = Nov 7, 1996 / 1245 days
Start of Mapping / Days since Start of Mapping = April 1, 1999 / 370 days Total Mapping Orbits = 4820
Total Orbits = 6423
The spacecraft continues to operate nominally in performing the beta supplement daily recording and transmission of science data. The mm026 sequence executed successfully from 00-90(3/30/00) through 00-93 (4/3/00). The mm027 sequence also successfully executed from 00-94 (4/4/00) through 00-96 (4/6/00). The mm028 sequence was successfully uplinked on 00-95 (4/5/00) and will begin execution on 00-97 (4/8/00). Included in the mm028 sequence, is the fifth and last star scan of the current MOC focus campaign, and a 24-hour period of MOLA off-nadir scans of the North and South Poles, scheduled on 00-99 through 00-100 (4/8-4/9/00). The mm029 sequence is scheduled to be uplinked on 00-99 (4/8/00) and will execute from 00-101 (4/10/00) through 00-103 (4/12/00).
The mz049 mini-sequence, containing the first four MOC star scans, has begun execution. The first three scans have completed to date. The MOC successfully acquired the desired stars from the first scan, but the stars were on the edge of their image. The second scan resulted in the MOC not seeing the stars. On the third scan the stars were acquired but were again on the edge of the image. The spacecraft attitude control errors were very small during each of the scans, leading us to believe there is some kind of knowledge error in the design of the scans, resulting in the stars not being centered in the image. AACS is currently analyzing the data and working with JPL and the MOC team to determine the cause of this error. Depending on the success of the fourth and fifth scans, additional scans may be performed in the next couple of weeks to ensure the MOC focus objective.
All subsystems are reporting nominal health.
There have been 14 uplinks to the spacecraft during the last week, including new star catalogs and ephemeris files, instrument command loads, and the mm028 and mm029 sequences. Total command files radiated to the spacecraft since launch is 4565.
Planning for the replacement of the monthly fixed-HGA periods for the purpose of obtaining radio science atmospheric occultation egress data, prohibited in normal beta supplement operations due to HGA gimbal constraints, proceeds. The first Radio Science occultation egress scans are scheduled for 5/20-5/22/00.